# QDepth-VLA: Quantized Depth Prediction as Auxiliary Supervision for Vision–Language–Action Models

### v1

@@ 1. Introduction
Large vision–language–action (VLA) models Black et al. (2024); Intelligence et al. (2025); Cheang et al. (2024, 2025) have recently emerged as a powerful paradigm for robotic learning. By grounding pre-trained vision–language models (VLMs) Beyer et al. (2024); Steiner et al. (2024); Wang et al. (2024) with action-generation capabilities, robots acquire strong generalization across diverse instructions and visual contexts. However, when applied to long-horizon or fine-grained manipulation tasks, these models often exhibit substantial performance degradation Yang et al. (2025a); Fan et al. (2025); Song et al. (2025); Zhu et al. (2025). The primary reason lies in a persistent gap between semantic understanding and geometric reasoning Goyal et al. (2024); Gervet et al. (2023).
@@
Without reliable 3D understanding, VLAs often misestimate object positions or gripper–object relations, leading to cascading errors during manipulation Song et al. (2025). Therefore, several recent works have explored incorporating geometric information into VLA models to enable a deeper understanding of the 3D physical environment. These approaches can be broadly generalized into three paradigms: direct 3D feature injection , 2D-projected 3D feature integration and auxiliary 3D information prediction. The first category injects encoded 3D representations, such as point clouds Li et al. (2025b) or depth maps Bhat et al. (2025), into the vision–language backbone or the action head. This strategy typically requires an additional encoder to process 3D features, increasing model complexity and computational cost. While providing explicit geometric cues, it may disrupt the powerful 2D priors learned during large-scale VLM pretraining, leading to degraded visual–language reasoning and understanding. The second category projects 3D features into 2D representations and feeds them into the VLM Li et al. (2025a). Although this preserves pretrained 2D priors, it inevitably introduces information loss in the projection process, which can hinder fine-grained manipulation performance. Compared to these two paradigms, enhancing geometric understanding through auxiliary visual prediction tasks, such as future depth maps estimation Zhang et al. (2025b), offers a more promising alternative. This approach not only preserves the strong 2D priors of pretrained VLMs, but also avoids the need for additional sensory inputs during inference, while encouraging the model to learn 3D-consistent spatial reasoning.
@@
However, existing works that employ depth-map–based visual prediction as auxiliary tasks Zhang et al. (2025b) have not achieved consistent performance improvements, and in some cases even indicate that introducing depth prediction as an auxiliary loss can be detrimental to policy learning due to noisy supervision and weak geometric grounding. The key challenges lie in three aspects. Firstly, the supervision quality of depth maps is often limited by insufficient spatial–temporal consistency across frames Yang et al. (2024); Chen et al. (2025), introducing substantial noise that weakens geometric grounding. Secondly, pixel-wise depth regression produces highly redundant learning signals, forcing the model to reconstruct every pixel rather than focusing on salient structural cues essential for manipulation. Thirdly, using a vision–language backbone to predict depth maps may interfere with its pre-trained semantic alignment, potentially degrading multimodal reasoning performance.
@@
    Comprehensive experiments on both the Simpler Li et al. (2024a) and LIBERO Liu et al. (2023) benchmarks demonstrate that QDepth-VLA substantially enhances policy performance, outperforming open $\pi_{0}$ Ren (2024) by 6.1% and 7.7% on average success rate, respectively. Moreover, QDepth-VLA achieves a 10.0% improvement in real-world robotic manipulation, validating its effectiveness and generalizability.
@@ 2. RELATED WORKS
3D spatial information has been widely explored to overcome the limitations of purely 2D-based models. Early efforts typically enhanced spatial perception by either lifting 2D inputs into 3D Gervet et al. (2023); Goyal et al. (2023, 2024) or directly fusing 2D visual features with 3D point clouds Yang et al. (2025b); Qu et al. (2025); Lin et al. (2025).
While these approaches demonstrate that incorporating 3D signals can significantly improve spatial perception and action precision, directly fusing 3D and 2D representations or relying solely on 3D features can disrupt the visual-language alignment established in large-scale VLM pretraining. To mitigate this, two alternative directions have been proposed: (1) Projecting 3D features into 2D space, as in BridgeVLA Li et al. (2025a), which renders 3D inputs into multi-view 2D images for compatibility with VLMs. (2) Independent 3D encoders encode geometric information for integration into the action head. This paradigm is employed by PointVLA Li et al. (2025b) and GeoVLA Sun et al. (2025), where specialized point cloud encoders supply 3D embeddings to modality-specific experts.
Despite these advances, point cloud reconstruction may lose fine-grained object details, and the modality gap between 2D RGB pretraining and 3D geometry remains a persistent challenge. By contrast, depth maps exhibit a much smaller gap with RGB images and thus offer a more natural bridge between 2D and 3D. Recent depth-based approaches have demonstrated this advantage. 3D-CAVLA Bhat et al. (2025) integrates Region of Interest (RoI) pooling with depth embeddings projected into VLM token space, achieving extraordinary multi-view performance, while 4D-VLA Zhang et al. (2025a) augments visual inputs with 3D coordinate embeddings to support both spatial alignment and temporal reasoning.
@@ 2.2. Auxiliary Visual Reasoning Tasks for VLA
A series of works focus on predicting future sub-goals, such as generating sub-goal images or short rollouts that visualize task progress. This strategy, as exemplified by CoT-VLA Zhao et al. (2025) , enhances temporal reasoning by conditioning action generation on both current and predicted states, but incurs high computational cost due to the difficulty of synthesizing realistic RGB predictions. Other researches Intelligence et al. (2025); Cheang et al. (2025) introduce object-centric signals, such as bounding boxes or spatial relations, which provide structured knowledge of entities and their interactions. More recently, latent future embeddings have been explored, where discrete action tokens predicted in a compressed latent space encode upcoming intentions. AgiBot World Colosseo AgiBot-World-Contributors et al. (2025) and UniVLA Bu et al. (2025) exemplify this paradigm, showing scalability through large-scale human video pretraining, yet such latent predictions often lack explicit 3D grounding and struggle to capture fine-grained geometry. Finally, some approaches turn to pixel-level 3D supervision, predicting dense depth or semantic maps to reinforce geometric awareness, as in 3D-VLA Zhen et al. (2024) and DreamVLA Zhang et al. (2025b). While sometimes effective for strengthening spatial reasoning, these signals are difficult to optimize directly and may overemphasize redundant low-level cues rather than the most relevant spatial structures.
@@ 3. Methodology
Figure 1. An overview of QDepth-VLA. (a) The overall architecture and training pipeline, where depth supervision is incorporated via a depth expert and latent prediction module. In co-training, the VQ-VAE van den Oord et al. (2017) encoder and codebook are frozen, while PaLI-Gemma 3B Beyer et al. (2024), the action expert, depth expert, SigLIP Zhai et al. (2023), and tokenizer are trainable. (b) The proposed hybrid attention mask, which integrates depth and visual tokens to enhance spatial reasoning and manipulation performance.
@@ 3.1. Depth Annotation
Since existing VLA datasets such as OXE dataset Padalkar et al. (2023) lack sufficient 3D annotations, we first generate monocular depth estimates for training. To ensure high-quality and spatial–temporal consistent depth sequences, we employ Video-Depth-Anything (ViDA) Chen et al. (2025), the current state-of-the-art monocular video depth estimation framework built upon a ViT-Large backbone, to acquire depth maps. Specifically, ViDA is applied to the main-view RGB frames from a subset of the OXE Padalkar et al. (2023) and LIBERO Liu et al. (2023) datasets to obtain temporally aligned relative depth annotations, providing reliable geometric supervision for depth tokenization and subsequent model training.
@@ 3.2. VQ-VAE Reconstruction
To represent depth compactly, we pretrain a Vector-Quantized Variational Autoencoder (VQ-VAE) van den Oord et al. (2017). Given a depth frame $\mathbf{x}$, the encoder $f_{\theta}(\cdot)$ produces a latent $\mathbf{z}_{e}=f_{\theta}(\mathbf{x})$, which is quantized to the nearest code vector in a codebook $\mathcal{C}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{K}\}$:
@@
We use $K=256$ codebook entries of dimension $d=160$, and train the VQ-VAE van den Oord et al. (2017) with the standard objective:
@@
where $\text{sg}[\cdot]$ denotes stop-gradient and $\beta=0.25$. In practice, we experiment with latent grid resolutions of $16\times 16$ and $32\times 32$. We find that the smaller $16\times 16$ configuration already achieves accurate depth reconstruction while remaining computationally efficient. The VQ-VAE van den Oord et al. (2017) is pretrained independently on each dataset using AdamW Loshchilov and Hutter (2017) with a learning rate of $1\times 10^{-5}$ to ensure stable convergence and reconstruction quality. The resulting pretrained model produces discretized depth code indices, which serve as supervisory targets for the depth expert in QDepth-VLA.
@@ 3.3. QDepth-VLA Architecture
QDepth-VLA adopts a unified and modular architecture built upon open $\pi_{0}$ Ren (2024), extending its VLA pipeline with an additional depth supervision branch. As shown in Fig. 1(a), the model consists of three parameterized modules: a pretrained vision–language model (VLM), an action expert, and a newly introduced depth expert. These modules are coordinated through a mixture-of-experts (MoE) structure and a carefully designed hybrid attention mask, enabling QDepth-VLA to jointly reason about geometry and control variants without disrupting pretrained representations.
@@
We choose PaliGemma-3B Beyer et al. (2024) as VLM backbone, which integrates SigLIP-based Zhai et al. (2023) vision encoding with Gemma’s Mesnard et al. (2024) language modeling capability. Input instructions are first tokenized using Gemma’s Mesnard et al. (2024) tokenizer, while the main-view RGB image is processed by the SigLIP Zhai et al. (2023) image encoder to obtain 256 visual tokens. These image tokens are concatenated with 20 text prefix tokens and fed into the Gemma Mesnard et al. (2024) decoder under full block attention to produce multimodal embeddings that capture both spatial and semantic cues. This pretrained VLM remains trainable during the training stage, allowing geometric adaptation to manipulation environment.
@@
The action expert is a transformer-based module responsible for translating multimodal embeddings and proprioceptive states into executable robot actions. It consists of stacked transformer layers with MLP-based encoders and decoders that integrate visual–language context from the VLM with proprioceptive features. This module, which is built upon the original open $\pi_{0}$ Ren (2024) action head, functions as the core control head of QDepth-VLA.
@@
To incorporate geometric reasoning, QDepth-VLA introduces a dedicated depth expert, architecturally aligned with the action expert (illustrated in Table 1). It takes the visual embeddings from the SigLIP encoder as input, before language fusion to avoid semantic interference. These embeddings are projected through a lightweight MLP, processed by a transformer backbone, and then passed to a shallow CNN decoder that predicts 256 depth tokens. Each predicted token corresponding to a latent vector is then aligned with the quantized tokens produced by the pretrained VQ-VAE van den Oord et al. (2017) encoder over its codebook. The pretrained VQ-VAE van den Oord et al. (2017) decoder is subsequently used to reconstruct the spatial depth map from these latent tokens when required. This discrete formulation enables QDepth-VLA to capture compact, structured geometric representations while maintaining optimization stability.
@@
As for hybrid attention mechanism, existing designs typically employ a standard causal attention structure, as seen in DreamVLA Zhang et al. (2025b) and CoT-VLA Zhao et al. (2025). However, since depth modalities inherently contain noise, directly fusing them under causal attention may introduce undesirable interference, potentially degrading action generation quality Zhang et al. (2025b). To address this issue, we redesign the hybrid attention mechanism (Fig. 1(b)) to more effectively regulate cross-modal information flow among text, image, depth, proprioception, and action tokens. To be specific:
@@ 3.4.1. Quantized Depth Supervision
During joint training, the depth expert predicts latent depth tokens, and these tokens are then used to compute logits with the encoded image features over the VQ-VAE van den Oord et al. (2017) codebook:
@@
where $i$ indexes latent spatial positions , $k$ indexes codebook entries ($K=256$) and $\tau$ represents temperature factor. A cross-entropy loss is applied using ground-truth code indices $z_{i}^{*}$ obtained from the pretrained VQ-VAE van den Oord et al. (2017):
@@ 3.4.2. Action Modeling
The Conditional Flow Matching (CFM) action loss Lipman et al. (2022) is identical to that of $\pi_{0}$ Black et al. (2024):
@@ 3.4.4. Optimization Setup
QDepth-VLA is trained using the AdamW Loshchilov and Hutter (2017) optimizer with decoupled weight decay. We set the learning rate for both the action expert and the VLM backbone to $5\times 10^{-5}$. A cosine learning rate scheduler with $200$ warm-up steps and a cycle length of $10^{7}$ steps is applied, ensuring stable optimization throughout training.
@@
Single-view VLA General VLA OpenVLA finetuned Kim et al. (2024) 84.7 88.4 79.2 53.7 76.5
CoT-VLA-7B Zhao et al. (2025) 87.5 91.6 87.6 69.0 81.1
Open $\pi_{0}$ Ren (2024) 77.2 84.0 83.6 66.0 77.7
3D-cloud-enhanced VLA SpatialVLA Qu et al. (2025) 88.2 89.9 78.6 55.5 78.1
Depth-enhanced VLA 3D-CAVLA Bhat et al. (2025) 86.1 94.7 82.9 66.8 82.6
@@
Multi-view VLA General VLA Diffusion Policy Chi et al. (2023) 78.3 92.5 68.3 50.5 72.4
Octo finetuned Team et al. (2024) 78.9 85.7 84.6 51.1 75.1
$\pi_{0}$-FAST finetuned Black et al. (2024) 96.4 96.8 88.6 60.2 85.5
$\pi_{0}$ finetuned Black et al. (2024) 96.8 98.8 95.8 85.2 94.2
UniVLA Bu et al. (2025) 96.5 96.8 95.6 92.0 95.2
3D-cloud-enhanced VLA GeoVLA Sun et al. (2025) 98.4 99.0 96.6 96.6 97.7
Depth-enhanced VLA 3D-CAVLA Bhat et al. (2025) 98.2 99.8 98.2 96.1 98.1
4D-VLA Zhang et al. (2025a) 88.9 95.2 90.9 79.1 88.6
DreamVLA Zhang et al. (2025b) 97.5 94.0 89.5 89.5 92.6

### v2

@@ 1. Introduction
Large vision–language–action (VLA) models Black20240AV; Intelligence202505AV; Cheang2024GR2AG; Cheang2025GR3TR have recently emerged as a powerful paradigm for robotic learning. By grounding pre-trained vision–language models (VLMs) Beyer2024PaliGemmaAV; Steiner2024PaliGemma2A; Wang2024Qwen2VLEV with action-generation capabilities, robots acquire strong generalization across diverse instructions and visual contexts. However, when applied to long-horizon or fine-grained manipulation tasks, these models often exhibit substantial performance degradation Yang2025RoboEnvisionAL; Fan2025LongVLAUL; Song2025ReconVLARV; Zhu2025BridgingVA. The primary reason lies in a persistent gap between semantic understanding and geometric reasoning Goyal2024RVT2LP; Gervet2023Act3D3F.
@@
Without reliable 3D understanding, VLAs often misestimate object positions or gripper–object relations, leading to cascading errors during manipulation Song2025ReconVLARV. Therefore, several recent works have explored incorporating geometric information into VLA models to enable a deeper understanding of the 3D physical environment. These approaches can be broadly generalized into three paradigms: direct 3D feature injection , 2D-projected 3D feature integration and auxiliary 3D information prediction. The first category injects encoded 3D representations, such as point clouds Li2025PointVLAIT or depth maps Bhat20253DCL, into the vision–language backbone or the action head. This strategy typically requires an additional encoder to process 3D features, increasing model complexity and computational cost. While providing explicit geometric cues, it may disrupt the powerful 2D priors learned during large-scale VLM pretraining, leading to degraded visual–language reasoning and understanding. The second category projects 3D features into 2D representations and feeds them into the VLM Li2025BridgeVLAIA. Although this preserves pretrained 2D priors, it inevitably introduces information loss in the projection process, which can hinder fine-grained manipulation performance. Compared to these two paradigms, enhancing geometric understanding through auxiliary visual prediction tasks, such as future depth maps estimation Zhang2025DreamVLAAV, offers a more promising alternative. This approach not only preserves the strong 2D priors of pretrained VLMs, but also avoids the need for additional sensory inputs during inference, while encouraging the model to learn 3D-consistent spatial reasoning.
@@
However, existing works that employ depth-map–based visual prediction as auxiliary tasks Zhang2025DreamVLAAV have not achieved consistent performance improvements, and in some cases even indicate that introducing depth prediction as an auxiliary loss can be detrimental to policy learning due to noisy supervision and weak geometric grounding. The key challenges lie in three aspects. Firstly, the supervision quality of depth maps is often limited by insufficient spatial–temporal consistency across frames Yang2024DepthAV; Chen2025VideoDA, introducing substantial noise that weakens geometric grounding. Secondly, pixel-wise depth regression produces highly redundant learning signals, forcing the model to reconstruct every pixel rather than focusing on salient structural cues essential for manipulation. Thirdly, using a vision–language backbone to predict depth maps may interfere with its pre-trained semantic alignment, potentially degrading multimodal reasoning performance.
@@
    Comprehensive experiments on both the Simpler Li2024EvaluatingRR and LIBERO Liu2023LIBEROBK benchmarks demonstrate that QDepth-VLA substantially enhances policy performance, outperforming open $\pi_{0}$ Ren2024OpenPiZero by 6.1% and 7.7% on average success rate, respectively. Moreover, QDepth-VLA achieves a 10.0% improvement in real-world robotic manipulation, validating its effectiveness and generalizability.
@@ 2. RELATED WORKS
3D spatial information has been widely explored to overcome the limitations of purely 2D-based models. Early efforts typically enhanced spatial perception by either lifting 2D inputs into 3D Gervet2023Act3D3F; Goyal2023RVTRV; Goyal2024RVT2LP or directly fusing 2D visual features with 3D point clouds Yang2025FP3A3; Qu2025SpatialVLAES; Lin2025Evo0VM.
While these approaches demonstrate that incorporating 3D signals can significantly improve spatial perception and action precision, directly fusing 3D and 2D representations or relying solely on 3D features can disrupt the visual-language alignment established in large-scale VLM pretraining. To mitigate this, two alternative directions have been proposed: (1) Projecting 3D features into 2D space, as in BridgeVLA Li2025BridgeVLAIA, which renders 3D inputs into multi-view 2D images for compatibility with VLMs. (2) Independent 3D encoders encode geometric information for integration into the action head. This paradigm is employed by PointVLA Li2025PointVLAIT and GeoVLA Sun2025GeoVLAE3, where specialized point cloud encoders supply 3D embeddings to modality-specific experts.
Despite these advances, point cloud reconstruction may lose fine-grained object details, and the modality gap between 2D RGB pretraining and 3D geometry remains a persistent challenge. By contrast, depth maps exhibit a much smaller gap with RGB images and thus offer a more natural bridge between 2D and 3D. Recent depth-based approaches have demonstrated this advantage. 3D-CAVLA Bhat20253DCL integrates Region of Interest (RoI) pooling with depth embeddings projected into VLM token space, achieving extraordinary multi-view performance, while 4D-VLA Zhang20254DVLASV augments visual inputs with 3D coordinate embeddings to support both spatial alignment and temporal reasoning.
@@ 2.2. Auxiliary Visual Reasoning Tasks for VLA
A series of works focus on predicting future sub-goals, such as generating sub-goal images or short rollouts that visualize task progress. This strategy, as exemplified by CoT-VLA Zhao2025CoTVLAVC , enhances temporal reasoning by conditioning action generation on both current and predicted states, but incurs high computational cost due to the difficulty of synthesizing realistic RGB predictions. Other researches Intelligence202505AV; Cheang2025GR3TR introduce object-centric signals, such as bounding boxes or spatial relations, which provide structured knowledge of entities and their interactions. More recently, latent future embeddings have been explored, where discrete action tokens predicted in a compressed latent space encode upcoming intentions. AgiBot World Colosseo AgiBotWorldContributors2025AgiBotWC and UniVLA Bu2025UniVLALT exemplify this paradigm, showing scalability through large-scale human video pretraining, yet such latent predictions often lack explicit 3D grounding and struggle to capture fine-grained geometry. Finally, some approaches turn to pixel-level 3D supervision, predicting dense depth or semantic maps to reinforce geometric awareness, as in 3D-VLA Zhen20243DVLAA3 and DreamVLA Zhang2025DreamVLAAV. While sometimes effective for strengthening spatial reasoning, these signals are difficult to optimize directly and may overemphasize redundant low-level cues rather than the most relevant spatial structures.
@@ 3. Methodology
Figure 1. An overview of QDepth-VLA. (a) The overall architecture and training pipeline, where depth supervision is incorporated via a depth expert and latent prediction module. In co-training, the VQ-VAE Oord2017NeuralDR encoder and codebook are frozen, while PaLI-Gemma 3B Beyer2024PaliGemmaAV, the action expert, depth expert, SigLIP Zhai2023SigmoidLF, and tokenizer are trainable. (b) The proposed hybrid attention mask, which integrates depth and visual tokens to enhance spatial reasoning and manipulation performance.
@@ 3.1. Depth Annotation
Since existing VLA datasets such as OXE dataset Padalkar2023OpenXR lack sufficient 3D annotations, we first generate monocular depth estimates for training. To ensure high-quality and spatial–temporal consistent depth sequences, we employ Video-Depth-Anything (ViDA) Chen2025VideoDA, the current state-of-the-art monocular video depth estimation framework built upon a ViT-Large backbone, to acquire depth maps. Specifically, ViDA is applied to the main-view RGB frames from a subset of the OXE Padalkar2023OpenXR and LIBERO Liu2023LIBEROBK datasets to obtain temporally aligned relative depth annotations, providing reliable geometric supervision for depth tokenization and subsequent model training.
@@ 3.2. VQ-VAE Reconstruction
To represent depth compactly, we pretrain a Vector-Quantized Variational Autoencoder (VQ-VAE) Oord2017NeuralDR. Given a depth frame $\mathbf{x}$, the encoder $f_{\theta}(\cdot)$ produces a latent $\mathbf{z}_{e}=f_{\theta}(\mathbf{x})$, which is quantized to the nearest code vector in a codebook $\mathcal{C}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{K}\}$:
@@
We use $K=256$ codebook entries of dimension $d=160$, and train the VQ-VAE Oord2017NeuralDR with the standard objective:
@@
where $\text{sg}[\cdot]$ denotes stop-gradient and $\beta=0.25$. In practice, we experiment with latent grid resolutions of $16\times 16$ and $32\times 32$. We find that the smaller $16\times 16$ configuration already achieves accurate depth reconstruction while remaining computationally efficient. The VQ-VAE Oord2017NeuralDR is pretrained independently on each dataset using AdamW Loshchilov2017DecoupledWD with a learning rate of $1\times 10^{-5}$ to ensure stable convergence and reconstruction quality. The resulting pretrained model produces discretized depth code indices, which serve as supervisory targets for the depth expert in QDepth-VLA.
@@ 3.3. QDepth-VLA Architecture
QDepth-VLA adopts a unified and modular architecture built upon open $\pi_{0}$ Ren2024OpenPiZero, extending its VLA pipeline with an additional depth supervision branch. As shown in Fig. 1(a), the model consists of three parameterized modules: a pretrained vision–language model (VLM), an action expert, and a newly introduced depth expert. These modules are coordinated through a mixture-of-experts (MoE) structure and a carefully designed hybrid attention mask, enabling QDepth-VLA to jointly reason about geometry and control variants without disrupting pretrained representations.
@@
We choose PaliGemma-3B Beyer2024PaliGemmaAV as VLM backbone, which integrates SigLIP-based Zhai2023SigmoidLF vision encoding with Gemma’s Mesnard2024GemmaOM language modeling capability. Input instructions are first tokenized using Gemma’s Mesnard2024GemmaOM tokenizer, while the main-view RGB image is processed by the SigLIP Zhai2023SigmoidLF image encoder to obtain 256 visual tokens. These image tokens are concatenated with 20 text prefix tokens and fed into the Gemma Mesnard2024GemmaOM decoder under full block attention to produce multimodal embeddings that capture both spatial and semantic cues. This pretrained VLM remains trainable during the training stage, allowing geometric adaptation to manipulation environment.
@@
The action expert is a transformer-based module responsible for translating multimodal embeddings and proprioceptive states into executable robot actions. It consists of stacked transformer layers with MLP-based encoders and decoders that integrate visual–language context from the VLM with proprioceptive features. This module, which is built upon the original open $\pi_{0}$ Ren2024OpenPiZero action head, functions as the core control head of QDepth-VLA.
@@
To incorporate geometric reasoning, QDepth-VLA introduces a dedicated depth expert, architecturally aligned with the action expert (illustrated in Table 1). It takes the visual embeddings from the SigLIP encoder as input, before language fusion to avoid semantic interference. These embeddings are projected through a lightweight MLP, processed by a transformer backbone, and then passed to a shallow CNN decoder that predicts 256 depth tokens. Each predicted token corresponding to a latent vector is then aligned with the quantized tokens produced by the pretrained VQ-VAE Oord2017NeuralDR encoder over its codebook. The pretrained VQ-VAE Oord2017NeuralDR decoder is subsequently used to reconstruct the spatial depth map from these latent tokens when required. This discrete formulation enables QDepth-VLA to capture compact, structured geometric representations while maintaining optimization stability.
@@
As for hybrid attention mechanism, existing designs typically employ a standard causal attention structure, as seen in DreamVLA Zhang2025DreamVLAAV and CoT-VLA Zhao2025CoTVLAVC. However, since depth modalities inherently contain noise, directly fusing them under causal attention may introduce undesirable interference, potentially degrading action generation quality Zhang2025DreamVLAAV. To address this issue, we redesign the hybrid attention mechanism (Fig. 1(b)) to more effectively regulate cross-modal information flow among text, image, depth, proprioception, and action tokens. To be specific:
@@ 3.4.1. Quantized Depth Supervision
During joint training, the depth expert predicts latent depth tokens, and these tokens are then used to compute logits with the encoded image features over the VQ-VAE Oord2017NeuralDR codebook:
@@
where $i$ indexes latent spatial positions , $k$ indexes codebook entries ($K=256$) and $\tau$ represents temperature factor. A cross-entropy loss is applied using ground-truth code indices $z_{i}^{*}$ obtained from the pretrained VQ-VAE Oord2017NeuralDR:
@@ 3.4.2. Action Modeling
The Conditional Flow Matching (CFM) action loss Lipman2022FlowMF is identical to that of $\pi_{0}$ Black20240AV:
@@ 3.4.4. Optimization Setup
QDepth-VLA is trained using the AdamW Loshchilov2017DecoupledWD optimizer with decoupled weight decay. We set the learning rate for both the action expert and the VLM backbone to $5\times 10^{-5}$. A cosine learning rate scheduler with $200$ warm-up steps and a cycle length of $10^{7}$ steps is applied, ensuring stable optimization throughout training.
@@
Single-view VLA General VLA OpenVLA finetuned Kim2024OpenVLAAO 84.7 88.4 79.2 53.7 76.5
CoT-VLA-7B Zhao2025CoTVLAVC 87.5 91.6 87.6 69.0 81.1
Open $\pi_{0}$ Ren2024OpenPiZero 77.2 84.0 83.6 66.0 77.7
3D-cloud-enhanced VLA SpatialVLA Qu2025SpatialVLAES 88.2 89.9 78.6 55.5 78.1
Depth-enhanced VLA 3D-CAVLA Bhat20253DCL 86.1 94.7 82.9 66.8 82.6
@@
Multi-view VLA General VLA Diffusion Policy Chi2023DiffusionPV 78.3 92.5 68.3 50.5 72.4
Octo finetuned Team2024OctoAO 78.9 85.7 84.6 51.1 75.1
$\pi_{0}$-FAST finetuned Black20240AV 96.4 96.8 88.6 60.2 85.5
$\pi_{0}$ finetuned Black20240AV 96.8 98.8 95.8 85.2 94.2
UniVLA Bu2025UniVLALT 96.5 96.8 95.6 92.0 95.2
3D-cloud-enhanced VLA GeoVLA Sun2025GeoVLAE3 98.4 99.0 96.6 96.6 97.7
Depth-enhanced VLA 3D-CAVLA Bhat20253DCL 98.2 99.8 98.2 96.1 98.1
4D-VLA Zhang20254DVLASV 88.9 95.2 90.9 79.1 88.6
DreamVLA Zhang2025DreamVLAAV 97.5 94.0 89.5 89.5 92.6